Two spherical balls of mass 10kg each are placed 10cm apart.find the gravitational force of attraction between them.

Grade:11

3 Answers

AAYUSH KESHAVA
11 years ago

Fgravitational=Gm1m2/r2

G=6.67300 × 10-11 m3 kg-1 s-2, m1=m2=10kg,r=0.1m

Fgravitational=(6.67300 × 10-11 m3 kg-1 s-2x10kgx10kg)/(0.1m)2

              =6.6730x10-7

Vyshnavi
7 years ago
Mass of each ball =m1=m2=10kgDistance between them =d=10cm=0.1mGravitational force of attraction =G=m1,m2/d2 F=(6.670*10-11)*10*10/(0.1)2 =6.67*10-11*102/6.01 =6.67*10-11*102/102 =6.67*10-11*102*102 =6.67*10-11*104 =
M Asad
5 years ago
When a 7000 N elevator moves from street level to the top of a building 300 m above the street level,what is the change in gravitational potential energy?